# Histopathology Slide Indexing and Search: Are We There Yet?

### v1

@@ Abstract
The search and retrieval of digital histopathology slides is an important task that has yet to be solved. In this case study, we investigate the clinical readiness of three state-of-the-art histopathology slide search engines, Yottixel, SISH, and RetCCL, on three patients with solid tumors. We provide a qualitative assessment of each model’s performance in providing retrieval results that are reliable and useful to pathologists. We found that all three image search engines fail to produce consistently reliable results and have difficulties in capturing granular and subtle features of malignancy, limiting their diagnostic accuracy. Based on our findings, we also propose a minimal set of requirements to further advance the development of accurate and reliable histopathology image search engines for successful clinical adoption.
@@ 1 Introduction
As histopathology slides become increasingly digitized, the process of manually searching and retrieving slides has become increasingly more time-consuming for pathologists [4, 8]. Recently, there has been growing interest in the development of automated search and retrieval systems for digital histopathology slides [5, 1, 16, 4, 6], which can help pathologists identify similar cases in developing a broader differential diagnosis. There has been widespread disagreements between the groups who have created these methods [12]. These systems leverage advances in artificial intelligence and machine learning to analyze large volumes of slides efficiently and accurately.
@@
Among the recent end-to-end systems proposed for histopathology image search, Yottixel [5], SISH [1], and RetCCL [16] have emerged as influential contenders, showcasing promising outcomes. Yottixel pioneered the processing of large-scale WSIs by introducing the concept of mosaics. Instead of extracting features from the entire WSI, Yottixel’s approach involves extracting features from mosaic tiles using a DenseNet-based feature extractor. Additionally, Yottixel incorporates the notion of barcoding [13, 14] to facilitate expedited retrieval by binarizing the extracted features. Similarly, SISH adopts a similar framework to Yottixel but incorporates an additional VQ-VAE-based [10] feature extractor. SISH also introduces advanced VEB tree-based [15] indexing and ranking algorithms to enhance the retrieval phase’s time complexity. In contrast, RetCCL employs the mosaic concept as well, but uniquely converts WSIs to mosaics after extracting features from tiled WSIs. Moreover, RetCCL introduces an effective contrastive-based feature extractor to improve feature quality.
@@
The introduction of successive systems claiming to have achieved state-of-the-art performance in the search and retrieval of histopathology slides, often supported by statistical metrics demonstrating agreement with trained pathologists’ judgments, raises the fundamental question of whether this problem has been satisfactorily addressed. Specifically, it prompts an inquiry into the readiness of these systems for deployment in clinical settings, where they can provide genuinely valuable information to pathologists, especially in challenging cases that even the most experienced pathologists struggle to reach a unanimous consensus. Given that the end users of these models are pathologists, their ability to extract pertinent and meaningful information from these systems becomes crucial in alleviating their workload burdens, regardless of the systems’ technical sophistication.
@@
In this case study, we evaluate these models on three previously unseen patient cases from our health system. Our objective is not to provide a quantitative statistical analysis of these models’ performance, as the respective authors have already accomplished that, but to offer a qualitative assessment of the usefulness of these models in the clinical setting. Our test cases encompass less controversial diagnoses, enabling a qualitative evaluation of each model’s performance under clinical conditions. To ensure fairness, we constructed each model’s database using a fixed number of slides from The Cancer Genome Atlas (TCGA) [17], while employing the same feature extractors as published by the original authors.
@@
In the subsequent sections, we provide an overview of our methods and approach towards implementation (Section 2). We then report our qualitative analysis of each model’s performance on three patient cases (Section 3). Finally, we discuss the current state of histopathology slide search engines and propose a set of minimal requirements for real-world deployment (Section 4).
@@ 2 Methods
Figure 1: A summary of indexing and retrieval methods of Yottixel [5] plus KimiaNet [11], SISH [1], and RetCCL [16]. (a) Feature extractor for Yottixel, (b) for SISH, and (c) for RetCCL.
@@ 2.1 Search Engines
Search engines commonly comprise two fundamental components: indexing and database generation, as well as ranking and retrieval. Given the large-scale nature of the images involved in this study, feature extraction becomes imperative for effective indexing. In terms of ranking, a suitable similarity measure is crucial, followed by post-processing steps to ensure result quality. In the supplementary methods section, we provide an overview of the feature extraction techniques, database indexing approaches, employed similarity measures, and result ranking methodologies utilized by the three primary methods under investigation (Fig. 1). Please be advised that all the hyper-parameters employed here are the parameters recommended by the authors of the models. We additionally share our code that we used to re implement methods that were not made available by the authors.
@@ 2.2 Database Slides
To ensure a fair comparison among all models, it was necessary to have consistent slides in the databases of each model. We constructed the database using slides available in TCGA [17]. Given our focus on lung, brain, and liver as primary sites for testing (see Section 2.3), it was essential to include slides from these sites in the databases. Additionally, to introduce a challenging aspect, slides from breast and colon were added to ensure that site retrieval experiments were not trivial. For each site, we randomly selected between 50 to 75 slides from subtypes containing at least 75 slides. The varying number of slides aimed to introduce class imbalance, mirroring real-world scenarios where some subtypes have more samples than others. Importantly, none of the slides in the database shared the same patient ID. The resulting database comprised 508 slides from 5 different sites and 8 different subtypes (Table 1).
It is worth noting that in each experiment, we utilized the pre-trained feature extractors provided by the respective authors. These feature extractors were trained on different datasets of varying sizes. The relatively small size of our database does not affect the performance of these models, as the only aspect influenced by data size is the feature extractor. As long as we have samples of the same class as the query within the database, a correctly functioning model should be capable of retrieving them.
Due to preprocessing criteria, we were not able to include 6 slides for Yottixel and 1 slide for SISH in the database. These slides are listed in Unprocessed Slides.
Table 1: Summary of database used for comparison experiments .Abbreviations are based on [6].
Primary Site Project Name (Subtype) Abbr. Num. Slides Num. Selected Slides
@@
Glioblastoma Multiforme
@@
2040
61
Brain Lower Grade Glioma
LGG
1543
69
Lymphoid Neasm Diffuse Large B-cell Lymphoma
DLBC
4
@@
Breast
Breast Invasive Carcinoma
BRCA
2704
72
Lymphoid Neasm Diffuse Large B-cell Lymphoma
DLBC
2
@@
Bronchus and lung
Lung Adenocarcinoma
LUAD
1359
68
Lung Squamous Cell Carcinoma
LUSC
1265
57
Mesothelioma
MESO
2
@@
Colon
Colon Adenocarcinoma
COAD
1307
59
Rectum Adenocarcinoma
READ
18
@@
Lymphoid Neasm Diffuse Large B-cell Lymphoma
DLBC
6
@@
Liver and intrahepatic bile ducts
Liver Hepatocellular Carcinoma
LIHC
778
72
Cholangiocarcinoma
CHOL
80
50
2.3 Tissue and Subtype Retrieval Experiments
We define “Tissue” as the tissue of cancer origin and “Subtype” as the final diagnosis, which is specific to the tissue type. All slides were sourced from real clinical cases at our institution to best approximate real-world scenarios. Team members who were responsible for algorithmic implementation were blinded from the ground truth to reduce the likelihood of bias.
For subtype experiments, we limited the search database to the slides with the same tissue type as the query. This was also true for patch retrieval experiments. A summary of the test slides used for these experiments is provided in Table 2.
Table 2: Summary of test slides and corresponding diagnoses used for comparision experiments.
Slide Subtype Tissue Source
Slide 1 Lung adenocarcinoma Lung Surgical resection
Slide 2 Astrocytoma/Low grade glioma Brain Surgical resection
Slide 3 Hepatocellular carcinoma Liver FNA Biopsy
3 Results
@@
3.1 Lung Adenocarcinoma Slide Retrieval
The highest performing model for retrieving Lung Adenocarcinoma (LUAD) slides using subtype search is reported as SISH, but RetCCL and Yottixel also show high performance on LUAD retrieval. In organ search, only RetCCL was able to retrieve slides of lung origin, while Yottixel and SISH performed poorly, but for subtype retrieval, all three methods were able to correctly retrieve several slides of LUAD (Figure 2a).
3.2 Low Grade Glioma Slide Retrieval
For Low Grade Glioma (LGG) case, SISH is reported as the highest performing algorithm for retrieval by subtype search, though RetCCL and Yottixel also have significant performances. Only RetCCL correctly retrieves LGG slides on subtype search whereas both Yottixel and SISH retrieve Glioblastoma cases, and all three methods correctly retrieve slides containing brain tissue for organ retrieval (Figure 2b).
3.3 Hepatocellular Carcinoma Slide Retrieval
For Hepatocellular Carcinoma (LIHC) subtype retrieval, SISH, Yottixel, and RetCCL have comparable performance, but for liver organ search, RetCCL shows higher performance. Both Yottixel and SISH performed poorly on liver tissue retrieval, but RetCCL achieves perfect performance in identifying liver slides, while for subtype search, Yottixel and SISH retrieve correct LIHC cases but RetCCL mistakenly retrieves several cases of Cholangiocarcinoma (Figure 2c).
3.4 Patch Retrieval on Lung Adenocarcinoma
For identifying similar patches of interest in Lung Adenocarcinoma, RetCCL is the most consistent at retrieving patches of lung glandular cell, but all three models may have difficulties in discriminating more granular, subtle features of malignancy. When querying a patch containing normal alveolar tissue, all three models correctly retrieve patches corresponding to normal alveolar tissue, but some patches from SISH and Yottixel include thickened layers of fibroblasts, suggesting potential limitations in their ability to discriminate subtle differences (Figure 3).

### v2

@@ Abstract
The search and retrieval of digital histopathology slides is an important task that has yet to be solved. In this case study, we investigate the clinical readiness of four state-of-the-art histopathology slide search engines, Yottixel, SISH, RetCCL, and HSHR on both unseen datasets and several patient cases. We provide a qualitative and quantitative assessment of each model’s performance in providing retrieval results that are reliable and useful to pathologists. We found high levels of performance across all models using conventional metrics for tissue and subtyping search. Upon testing the models on real patient cases, we found the results were still less than ideal for clinical use. Based on our findings, we propose a minimal set of requirements to further advance the development of accurate and reliable histopathology image search engines for successful clinical adoption.
@@ 1 Introduction
Figure 1: A summary of feature extraction and database creation processes proposed by (a) Yottixel [12], (b) SISH [1], (c) RetCCL [24], and (d) HSHR [14]. The feature extractor of Yottixel is switched with KimiaNet [18].
As histopathology slides become increasingly digitized, the process of manually searching and retrieving slides has become increasingly more time-consuming for pathologists [10, 15]. Recently, there has been growing interest in the development of automated search and retrieval systems for digital histopathology slides [12, 1, 24, 10, 11, 14], which can help pathologists identify similar cases in both developing and narrowing a differential diagnosis. These systems leverage advances in artificial intelligence and machine learning to analyze large volumes of slides efficiently and accurately.
@@
Among the recent end-to-end systems proposed for histopathology image search, Yottixel [12], SISH [1], RetCCL [24], and HSHR [14] have emerged as influential contenders, showcasing promising outcomes. Yottixel pioneered the processing of large-scale WSIs by introducing the concept of mosaics. Instead of extracting features from the entire WSI, Yottixel’s approach involves extracting features from mosaic tiles using a DenseNet-based feature extractor. Additionally, Yottixel incorporates the notion of barcoding [20, 21] to facilitate expedited retrieval by binarizing the extracted features. Similarly, SISH adopts a framework very similar to Yottixel but incorporates an additional VQ-VAE-based [16] feature extractor. SISH also introduces advanced VEB tree-based [22] indexing and ranking algorithms to enhance the quality of the retrieved samples. In contrast, RetCCL employs the mosaic concept as well, but uniquely converts WSIs to mosaics after extracting features from tiled WSIs. Moreover, RetCCL introduces an effective contrastive-based feature extractor to improve feature quality. Finally, HSHR expands the idea of using Self-Supervised Learning (SSL) to both extract the mosaics and creating hash codes from them. It uses SimCLR [2] to train the feature extractor and uses MOCO [9] to train the Cluster-Attention Hash Encoder (CaEncoder). The results are then processed to create a hypergraph which leads to similarity-based WSI retrieval.
@@
The introduction of successive systems claiming to have achieved state-of-the-art performance in the search and retrieval of histopathology slides, often supported by statistical metrics demonstrating agreement with trained pathologists’ judgments, raises the fundamental question of whether this problem has been satisfactorily addressed. Specifically, it prompts an inquiry into the readiness of these systems for deployment in clinical settings, where they can provide genuinely valuable information to pathologists, especially in challenging cases where even the most experienced group of pathologists struggle to reach a consensus.
@@
In this case study, we evaluate these models on patient cases from our health system and several external datasets. Our objective is to provide a quantitative analysis of these models’ performances on unseen slides while offering a qualitative assessment of the usefulness of these models in the clinical setting and potential areas of improvement. To ensure fairness, we constructed each model’s database using a fixed number of slides from The Cancer Genome Atlas (TCGA) [25], while employing the same feature extractors as published by the original authors.
@@
In subsequent sections, we provide an overview of our methods and approach towards implementation (Section 2). We then report our quantitative and and qualitative analysis of model performance (Section 3). Finally, we discuss the current state of histopathology slide search engines and propose a set of minimal requirements for real-world deployment based on our findings (Section 4).
@@ 2.1 Search Engines
Search engines commonly comprise two fundamental components: indexing and database generation, as well as ranking and retrieval. Given the large-scale nature of the images involved in this study, feature extraction becomes imperative for effective indexing. In terms of ranking, a suitable similarity measure is crucial, followed by post-processing steps to ensure result quality. In the supplementary methods section, we provide an overview of the feature extraction techniques, database indexing approaches, employed similarity measures, and result ranking methodologies utilized by the four primary methods under investigation (Fig. 1). Please be advised that all the hyper-parameters employed here are the parameters recommended by the authors of the models. We additionally share our code that we used to re implement methods that were not made available by the authors.
@@
Finally, HSHR first trains a encoder in a self-supervised manner on a small subset of patches extracted from database slides. By clustering the features from this encoder, it creates mosaics for each slide, and then passes the the features of mosaics to the CaEncoder. By following teh guidelines of MOCO [9], they train the desired CaEncoders and this way, they are able to create hashings and weights for each slides. Hashing and weights are then incorporated into building a hypergraph for the database. Every query slide is then considered a new node and hyperedge in this hypergraph database and similarity scores can be calculated for it. Per the authors’ emphasize on global perception of WSIs, HSHR is not designed to be used with patch retrieval tasks.
A more detailed explanation of these models also can be found in Supplementary Materials (Supplementary Section Search Engines Methods). Specifically, Supplementary Algorithms 1 to 4 would summarize the process a query slide would undergo in all the discussed methods. Moreover, time complexity of various stages of ranking and retrieval of these methods are also juxtaposed in Complexity Analysis.
@@ 2.2 Database Slides
To ensure a fair comparison among all models, it was necessary to have consistent slides in the databases of each model. We constructed the database using slides available in TCGA [25]. Given our focus on lung, brain, and liver as primary sites for testing (see Section 2.4), it was essential to include slides from these sites in the databases. Additionally, to introduce a challenging aspect, slides from breast and colon were added to ensure that site retrieval experiments were not trivial. For each site, we randomly selected between 50 to 75 slides from subtypes containing at least 75 slides. The varying number of slides aimed to introduce class imbalance, mirroring real-world scenarios where some subtypes have more samples than others. Importantly, none of the slides in the database shared the same patient ID. The resulting database comprised 508 slides from 5 different sites and 8 different subtypes (Datasets).
It is worth noting that in each experiment, we utilized the pre-trained feature extractors provided by the respective authors (except for the backbone of HSHR, see Supplementary Section Search Engines Methods). These feature extractors were trained on different datasets of varying sizes. The relatively small size of our database does not affect the performance of these models, as the only aspect influenced by data size is the feature extractor. As long as we have samples of the same class as the query within the database, a correctly functioning model should be capable of retrieving them.
Due to preprocessing criteria, we were not able to include 6 slides for Yottixel, 1 slide for SISH,. 4 slides for HSHR in the database. These slides are listed in Datasets.
2.3 Test Datasets
In order to conduct fair quantitative experiments and avoid data leakage, we needed to acquire test slides that were not seen by the encoders of these models. Table 1 summarizes the all the datasets used in validation experiments. Except from the in house UCLA dataset, all other datasets are downloaded from the publicly available Cancer Imaging Archive database [3] (Datasets). For all the datasets, we made sure not to have samples from the same patient using the patient identifiers provided.
All UCLA slides were sourced from real clinical cases at our institution to best approximate real-world scenarios. Team members who were responsible for algorithmic implementation were blinded from the ground truth to reduce the likelihood of bias.
Table 1: Summary of test slides used for experiments. Abbreviations are based on [11].
Experiment Slides Dataset Site Diagnosis
UCLA
slide1
In House
lung
LUAD
slide2
In House
@@ 2.3 Test Datasets
LGG
slide3
In House
liver
LIHC
Reader Study
MSB-09151-01-11
CMB-CRC
colon
COAD
MSB-09977-01-22
CMB-LCA
lung
LUSC
Her2Pos_Case_66
Yale Her2+ Cohort
breast
BRCA
Microscope Study
34 slides
CPTAC-GBM (Leica)
brain
@@
34 slides
UPENN-GBM (Hamamatsu)
brain
GBM
HER2+ prediction
93 slides
Yale Her2+ Cohort
breast
BRCA
97 slides
Yale Her2- Cohort
breast
BRCA
Ablation
85 slides
Yale Trastuzumab Cohort
breast
BRCA
2.4 Experiments
In general, we have three types of experiments: site (tissue) retrieval, subtype retrieval, and patch retrieval. We define “site” as the tissue of cancer origin and “subtype” as the final diagnosis, which is specific to the tissue type. For patch retrieval tasks, the models should return the closest patches to a query patch as opposed to WSIs to WSIs. For subtype and patch retrieval experimetns, we limited the search database to the slides with the same tissue type as the query.
Consistent with prior work on WSI search algorithms, we chose majority top-k accuracy (mMV@K) and mean average precision (mAP@K) as our quantitative metric of performance, which returns the predicted label by majority vote amongst the top K slides. For tissue search, K has commonly been 10 for mMV and 5 for mAP. As with prior studies, we choose mMV@1,3,5 and mAP@3,5 for subtype search (See Supplementary Algorithms 5 and 6).
We have designed 5 experiments to validate the methods for different purposes. The UCLA experiment aims to bring qualitative evaluation to the 3 in house slides by evaluating them for all three tasks. Reader study is designed to bring pathologists’ point of view to the quality retrieved patches by the models. The microscope study is intended to measure the robustness of the models’ performances with respect to different microscope brands. The Her2+ prediction tries to answer the question whether there is evidence that the models would perform differently given different sub-subtype. And finally, since the authors’ of Yottixel mentioned their model would perform better using KimiaNet instead of DenseNet pretrained on ImageNet, we designed the ablation study to measure this change in performance.
3 Results
Table 2: Evaluation results of different methods on UCLA slides for primary site retrieval task.
Method UCLA Slides MV@1 MV@3 MV@5 MV@10 AP@3 AP@5
YOTTIXEL + KimiaNet Slide1
@@ 3 Results
Slide2
1
1
1
1
1
1
Slide3
0
0
0
0
0
0
SISH + DenseNet Slide1
0
0
0
Slide2
1
1
1
1
1
0.888
Slide3
0
0
0
0
RetCCL Slide1
1
0
1
1
1
0.750
Slide2
0
1
1
1
0.583
0.679
Slide3
1
1
1
1
1
1
HSHR Slide1
1
1
1
1
1
1
Slide2
1
1
1
1
1
1
Slide3
0
0
0
0
0.500
0.500
Table 3: Evaluation results of different methods on UCLA slides for subtype retrieval task.
Method UCLA Slides MV@1 MV@3 MV@5 AP@3 AP@5
YOTTIXEL + KimiaNet Slide1
0
0
1
0.500
0.533
Slide2
0
0
0
0
0
Slide3
1
1
1
1
1
SISH + DenseNet Slide1
1
1
1
Slide2
0
0
0
0
0
Slide3
1
1
1
1
1
RetCCL Slide1
1
1
1
1
1
Slide2
1
1
1
1
1
Slide3
1
0
0
1
0.700
HSHR Slide1
1
1
1
1
0.867
Slide2
1
0
0
1
1
Slide3
0
0
0
0.500
0.500
@@
3.1 Tissue and subtype retrieval
A quantitative analysis of model performance on our patient cases is reported in Table 2 and Table 3. While we cannot arrive at definitive conclusions on algorithmic performance due to the few number of cases here, we observed higher mMV and mAP metrics for RetCCL on both tissue and subtype retrieval relative to the other models.
Extended Results and Extended Results show the results of our quantitative evaluations on GBM and BRCA datasets encompassing 344 unseen slides. Of note, UPENN GBM was not feasible for RetCCL due to extensive computational burdens. When testing the models on brain tissue retrieval, we found that RetCCL had the highest performance at mMV @ 10 although there is almost a 15 point drop in performance versus the mMV @ 10 score reported by the authors at 90.21. For subtyping on GBM versus LGG, we see that HSHR and SISH are the top performers at mMV@5 for GBM versus LBB subtyping, which are lower than prior work by the authors of HSHR showing a mMV@5 of 0.937 and 0.916 for HSHR and SISH, respectively, when tested on 3580 TGCA GBM and LGG slides.
For breast tissue retrieval on the Yale Trastuzumab dataset, Yotixxel performed the best on both metrics with a mMV@10 of 0.588 and mAP@5 of 0.650. Of note, the authors of RetCCL previously showed that Yotixxel achieved a mMV@10 of 0.663 relative to RetCCL’s score of 0.914 on a set of frozen WSIs. Subtype search on the BRCA dataset was not performed as all of the cases were of the same diagnosis.
3.2 Visual review of query results
To better investigate discrepancies in performance, we reviewed the top five ranked results on subtyping and tissue search for three different WSI slides from our own patient cases as illustrated in Figure 2. Additional details on patient cases and slide preparation methods can be found in the Supplementary Section on Patient Cases.
In Figure 2, we see several errors made on tissue and subtype search. We also review of patch-level results for Yotixxel, SISH, and RetCCL on two patches from our LUAD case, one showcasing tumoral tissue and the other with normal alveolar tissue (Figure 3). HSHR was excluded due to its inability for patch-level search. We found that all three algorithms are capable of retrieving patches containing tumoral and alveolar tissue but there were visual discrepancies in some of more granular features on slides. For example, while all models retrieve patches corresponding to alveoli, all three models return patches with varying degrees of necrosis, inflammation, and hyperplasia, which may point pathologists towards different diagnoses and treatments.
3.3 Reader study
We next compared model performance on patch-level retrieval results qualitatively. Supplementary Figure 1 shows the Mean Opinion Score (MOS) of seven pathologists on the top three ranked results when querying a patch containing tumor from three WSI H&E slides. We include Yottixel, RetCCL, and SISH but not HSHR given the latter’s inability to perform patch-level search. For consistency, we used MOS as an evaluation metric based on prior studies on the quality of WSI search results. We found SISH had overlapping performance with Yotixxel and RetCCL due to higher variance but when comparing RetCCL versus Yotixxel, we see a statistically significant improvement in RetCCL performance. However, the Fleiss’ Kappa for all algorithms combined was 0.131 suggesting low rates of agreement amongst pathologists. Quantitative performance metrics are provided in Supplementary Tables 5 and 6 on our reader study slides.
3.4 HER2+ prediction
To test the richness of feature representations, we compared the ability of the four models in distinguishing between HER2+ and HER2- BRCA (Supplementary Figures 3,4,5, & 6). At present, immunohistochemistry is required for the interpretation of HER2 status although recent work using a CNN-based architecture was capable of predicting HER2 positivity with an AUC of 0.81 on untested datasets ([4]). All models tended to predict HER2- slides with greater precision despite the nearly 50-50 distribution of HER2- and HER2+ cases in the dataset, suggesting an ability to learn subtle features specific to HER2 status.
For this experiment, we calculated MV@10 and AP@5 for tissue retrieval task in both Her2+ and Her2- cohorts. Then using the Shapiro-Wilk normality test, and Levene’s homogeneity of variances test, we checked the distribution of these two metrics. Since they did not pass the Independent T test requirements, we conducted a non-parametric Mann-Whitney U Test. We concluded that all 4 models would have a better performance in terms of AP@5 on Her2- cohort for the site retrieval task. However, for MV@10, the evidence was only significant for Yottixel and HSHR (Reader study).
3.5 Ablation study
Figure Reader study shows comparative analysis of Yottixel and SISH algorithms using two distinct networks, Kima Net and Densenet, on the Yale Trastuzumab dataset is presented. The figure is subdivided into two subplots, each illustrating the performance of a base algorithm, Yottixel or SISH, evaluated across six different metrics: mmV @ 1, 3, 5, 10 and mAP @ 3 and 5. Configurations employing Kima Net generally outperform those using Densenet across our evaluation metrics. Disparities in performance underscore the impact of the network choice on the efficacy of the algorithms, highlighting the importance of optimal network-algorithm pairing.
3.6 Microscope study
We investigated differences in performance for Yotixxel, SISH, and HSHR on two GBM datasets utilizing different microscopes, UPENN and CPTAC GBM, to test model generalizability. As RetCCL was not compatible with the UPENN dataset, it was excluded from this experiment. We found that all three algorithms had overlapping results with p-values of $>$ 0.05 across both subtype and tissue retrieval at mMV @ 5 and 10 on both GBM datasets, indicating that differences in microscopes did not affect performance. Full results can be found in the Reader study.